-54=3(2+5m) I really have no idea.
Answer by stanbon(75887)

54=3(2+5m)
-----------
Divide both sides by 3 to get:
18 = 2+5m
----
16 = 5m
---
m = 16/5 = 3 1/5
===========================
Cheers,
Stan H.